Covered Interest Arbitrage Assume the following information: Given this information, is covered interest arbitrage worthwhile for Mexican investors who have pesos to invest? Explain your answer.

            A7DZW6LkDmqBAAAAAElFTkSuQmCC
Answer #1

Yes, covered interest arbitrage is worthwhile for Mexican investors who have pesos to invest. This is because an arbitrage profit can be earned.

A Mexican investor with 1000 pesos (for example) can earn an arbitrage profit with these steps :

  • Borrow 1000 pesos. Convert into $ at the spot exchange rate. $ received = 1,000 * 0.100 = $100
  • Invest $100 at the US interest rate for 180 days. $ received after 180 days = $100 * (1 + (2.5% * (180 / 360))) = $102.50
  • Enter into a forward contract today to convert $102.50 into pesos after 180 days at the forward exchange rate.
  • After 180 days, convert $ into pesos using the forward contract. Pesos received = $102.50 / 0.098 = 1,045.92 pesos
  • Repay the original 1,000 pesos borrowed with interest. Pesos to repay = 1,000 * (1 + (3% * (180 / 360))) = 1,030 pesos
  • Arbitrage profit = pesos received after 180 days - pesos to repay after 180 days
  • Arbitrage profit = 1,045.92 - 1,030 = 15.92 pesos